Selected information from Consolidated Income Statement, year ended 31 December 2018 (USD millions);

Gross profit 7,200
Other operating expenses 2,900
Operating profit 4,300
Interest income 100
Interest expense 300
Income before taxes 4,100
Income taxes 1,400
Net income 2,700
Other
Depreciation and amortization
500


The EBITDA interest coverage ratio is closest to:

16.00x.

9.33x.

14.33x.
Answer #1

EBITDA= earnings before interest, taxes, depreciation and amortization

EBIT= operating profit

For EBITDA, just we need to add only other depreciation and amortization

Other depreciation and amortization= 500

EBITDA= 4300+ 500= 4800

EBITDA interest coverage ratio= EBITDA/ interest Expense

Interest Expense=300

EBITDA interest coverage ratio= 4800/300=16.00x